Prove the theorem

The elass of regular languages is closed under union operation. In othe words, if Al

and A2 are regular languages, so is (Ai UA2)"for the following languageL-twiw acepts even mumber of 'a' or even number of "b'.


1
Expert's answer
2021-04-02T11:53:39-0400

Prove that for regular languages "L_1" and "L_2" that "L_1\lor L_2" is regular.

Let "M_1=(S_1,\sum, \delta_1,s_1,F_1)" and "M_2=(S_2,\sum, \delta_2,s_2,F_2)" be DFA for "L_1" and "L_2"

DFA Construction (the Product Construction):

We claim the DFA "M=(S_1\times S_2,\sum, \delta,(s_1,s_2),F)" where

"\delta((s_1,s_2),a)=(\delta_1(s_1,a),\delta_2(s_2,a))" for all "s_1\isin S_1,s_2\isin S_2" and "a\isin \sum" ,

"F=(F_1\times S_2)\lor (S_1\times F_2)"

accepts "L_1\lor L_2"

i.e. "L(M)=L(M_1)\lor L(M_2)" .

Proof of correctness: For every string "w" , we have:

"\hat{\delta}((s_1,s_2),w)=(\hat{\delta}_1(s_1,w),\hat{\delta}_2(s_2,w))"

"\hat{\delta}((s_1,s_2),w)\isin F" iff "\hat{\delta}_1(s_1,w)\isin F" or "\hat{\delta}_2(s_2,w)\isin F" .

"L_1\lor L_2" is regular since there is a DFA accepting this language.